If a = 2i + 4j and b = 6j-4k get the smallest angle between the two vectors by using:
(a) Cross Product
(b) Dot Product
Answer #1

a) we know axb=1ballbl sino. lal, 1bl are magnitude of vector a and b. If azax, itfai+2, K and b= xzityzj + z k then lal=V x²axb - -16i +8 5+12 K. labl= √(-16)+8) 2+ (2) ► - 1464 sino = 656 464 152 *V20 41.91° Q = S ( 65 ) (52+520) a,b zla) 161 cosce